# Equivariant plug-and-play image reconstruction

### v1

@@ 5.4 Sampling from RED prior
The gain in stability provided by the proposed approach opens the door to more robust sampling algorithms relying on implicit denoising priors, such as ULA, where a sufficiently large number of iterations is required in order to obtain good estimators. We show in Figure 8 the estimated mean and variance obtained with the ULA algorithm, with both equivariant and non-equivariant DRUNet backbone denoisers. In the non-equivariant setting, we observe similar artifacts to those obtained in the deterministic case on both the estimated mean and variance; these artifacts vanish in the equivariant case.
The plots of Figure 8 further illustrates this phenomenon. We notice that after a few hundred iterations, the reconstruction quality with the non-equivariant ULA algorithm collapses, yielding an irrelevant MCMC chain. While enforcing equivariance improves the situation, increasing the noise level in the denoiser in PnP can significantly improve the situation, as shown on the rightmost plot of Figure 8.
@@
Figure 8: Sampling with the ULA algorithm on a motion deblurring and the DRUNet prior, in the standard and equivariant cases. Left: PSNR of a sample from the MCMC chain for a sample of the Set3C dataset. Right: average PSNR over the Set3C dataset as a function of $\sigma$ in the denoising prior.
@@ 6 Limitations
While our experiments show that the proposed method enables to always improve both the stability of the algorithmic scheme and the reconstructed image quality over its non-equivariant counterpart, we stress that it remains prone to divergence and/or hallucinating artifacts. This can for example be seen in Figure 4 in the case of the DRUNet backbone for example, which shows decreasing mean PSNR after a certain number of iterations, or in Figure 5 for the SCUNet backbone, which shows important reconstruction artifacts. In particular, the proposed equivariant approach when using a SCUNet backbone architecture does not clearly improve over its non-equivariant counterpart.
@@ 7 Conclusion
In this work, we have proposed a simple yet efficient method for ensuring approximate equivariance of implicit denoising priors. In essence, the method amounts to sampling and applying a group action at random at each step of the algorithm. In spite of its simplicity, this method shows interesting theoretical properties. In the case of linear denoisers, for example, it allows us to enforce the symmetry of the Jacobian, which is a cornerstone property when aiming to derive explicit priors from implicit denoising priors. Furthermore, it can be shown that the Lipschitz constant of the equivariant linear denoiser can only be lower than that of its non-equivariant counterpart, thus improving the stability of the resulting PnP algorithm. We showcase the symmetrization effect of equivariance on the Jacobian of several architectures, and illustrate its stabilization effect for three families of imaging algorithms, namely PnP, RED, and ULA. Importantly, this stabilization procedure is not detrimental to the reconstruction quality, as often observed in the literature. However, despite the significant improvements brought by equivariance both in terms of stability and image reconstruction quality, the proposed method remains prone to divergence and artifact contamination in the reconstructions.
@@ Appendix A Details on the non-linear example
Appendix B Equivariant algorithms

### v2

@@ 5.4 Sampling from RED prior
The gain in stability provided by the proposed approach opens the door to more robust sampling algorithms relying on implicit denoising priors, such as 3, where a sufficiently large number of iterations is required in order to obtain good estimators. We show in Figure 9 the estimated mean and variance obtained with the 3 algorithm, with both equivariant and non-equivariant DRUNet backbone denoisers. In the non-equivariant setting, we observe similar artifacts to those obtained in the deterministic case on both the estimated mean and variance; these artifacts vanish in the equivariant case.
@@
Figure 9: Sampling posterior mean and variance of a motion deblurring problem on a BSD10 sample for a DRUNet backbone plugged in the 3 algorithm.
The leftmost plot of Figure 10 further illustrates this phenomenon. We notice that after a few hundred iterations, the reconstruction quality with the non-equivariant 3 algorithm collapses, yielding an irrelevant MCMC chain. While enforcing equivariance improves the situation, increasing the noise level in the denoiser in PnP can significantly improve the situation, see the rightmost plot of Figure 10.
Figure 10: Sampling with the 3 algorithm on a motion deblurring and the DRUNet prior, in the standard and equivariant cases. Left: PSNR of a sample from the MCMC chain for a sample of the Set3C dataset. Right: average PSNR over the Set3C dataset as a function of $\sigma$ in the denoising prior.
@@ 6 Limitations
While our experiments show that the proposed method improves both the stability of the algorithmic scheme and the reconstructed image quality over its non-equivariant counterpart, we stress that it remains prone to divergence and/or hallucinating artifacts. This can for example be seen in Figure 4 in the case of the DRUNet backbone, which shows decreasing mean PSNR after a certain number of iterations, or in Figure 5 for the SCUNet backbone, which shows important reconstruction artifacts. In particular, the proposed equivariant approach does not clearly improve over its non-equivariant counterpart when using a SCUNet backbone.
@@ 7 Conclusion
In this work, we proposed a simple yet efficient method for ensuring approximate equivariance of implicit denoising priors. In essence, the method amounts to sampling and applying a group action at random at each step of the algorithm. In spite of its simplicity, this method shows interesting theoretical properties. In the case of linear denoisers, for example, it allows us to enforce the symmetry of the Jacobian, which is a cornerstone property when aiming to derive explicit priors from implicit denoising priors. Furthermore, it can be shown that the Lipschitz constant of the equivariant linear denoiser can only be lower than that of its non-equivariant counterpart, thus improving the stability of the resulting PnP algorithm. We showcase the symmetrization effect of equivariance on the Jacobian of several architectures, and illustrate its stabilization effect for three families of imaging algorithms, namely PnP, RED, and ULA. Importantly, this stabilization procedure is not detrimental to the reconstruction quality, as often observed in the literature. However, despite the significant improvements brought by equivariance in terms of stability and image reconstruction quality, the proposed method remains prone to divergence and artifact contamination in reconstructions.
Acknowledgements
This work was supported by the BrAIN grant (ANR-20-CHIA-0016) and was granted access to the HPC resources of IDRIS under the allocation 2023-AD011014344 made by GENCI. J. T. is supported by the ANR’s JCJC grant UNLIP.
@@ Appendix A Details on the non-linear example
Appendix B MC sampling and Reynolds averaging
We compare in Table 5 the performance of the equivariant architecture when training with the proposed Monte-Carlo (MC) scheme vs the true averaging. It shows no difference in final performance while the MC strategy decrease the computational complexity by a factor 4.
Architecture Dataset Monte-Carlo Sample Reynolds Average
DnCNN BSD10 $30.698\pm 1.645$ $30.684\pm 1.645$
DRUNet fastMRI $30.678\pm 0.740$ $30.646\pm 0.752$
LipDnCNN Set3C $32.705\pm 0.868$ $32.706\pm 0.868$
Table 5: Performance of algorithms from Fig. 4 when relying on Monte-Carlo estimates and averaged equivariant architectures.
Appendix C Equivariant algorithms